If the weights of 600 students are normally distributed with a mean of 50 kilograms and a variance of 16 kg




b.) How many students with weights between 50 kgs and 55 kgs. (5 points)




1
Expert's answer
2022-05-16T23:55:36-0400

Here the everage weights of 600 students is "50" kg with standard deviation "\sigma=\sqrt{variance}=4"

Let "X" be a random variable denotes the weight of the students.

Then "X" is normally distributed with mean "50" kg and standard deviation "4" kg.

Therefore we have "\mu =50" and "\sigma=4"  .

Let us take "Z= \frac{X-\mu}{\sigma}" . Then "Z=\frac{X-50}{4}".

Now we have to find "P(50<X<55)."

"P(50<X<55)=P(\frac{50-50}{4}<Z<\frac{55-50}{4})"

"=P(0<Z<1.25)"

"=0.8944-0.5=0.3944"

So the number of students with weights between "50" kg and "55" kg is "=600*0.3944)=236.64\approx237"